# Transversal Hamilton cycles in digraph collections

### v1

@@ 5 Extremal case
For vertices in $X\setminus X^{\prime}$, we consider the following claim. Notice that the position (in $A$ or $B$) of vertices in $X\setminus X^{\prime}$ does not affect our proof, so if $v\in X$ is a vertex moved from $B$ to $A$, then we only need consider $v\in A$.
Claim 5.4.
Assume $X\setminus V(P^{2})=\{x_{1},\ldots,x_{s}\}$ with $s\leq\sqrt{\epsilon}n$. If $x_{i}\in X\cap Y$, then there exist $c_{i}^{1},c_{i}^{2},c_{i}^{3},c_{i}^{4}\in\mathcal{C}_{2}\setminus({\rm col}(P^{2})\cup\{c_{\ell}^{1},c_{\ell}^{2},c_{\ell}^{3},c_{\ell}^{4}:{\ell}\in[i-1]\})$ and $x_{i}^{1},\,x_{i}^{2},\,x_{i}^{3},x_{i}^{4}\in Z\setminus(V(P^{2})\cup\{x_{\ell}^{1},\,x_{\ell}^{2},\,x_{\ell}^{3},\,x_{\ell}^{4}:{\ell}\in[i-1]\})$ such that ${x_{i}^{k}x_{i}}\in E(D_{c_{i}^{k}})$ for $k\in[2]$ and ${x_{i}x_{i}^{k}}\in E(D_{c_{i}^{k}})$ for $k\in\{3,4\}$.
Proof of Claim 5.4.
Suppose there exists an $i_{0}\in[s]$ such that Claim 5.4 holds for all $i\in[i_{0}-1]$ but does not hold for $i_{0}$. Assume for convenience that $x_{i_{0}}\in X\cap A$. Denote $\mathcal{C}_{2}^{\prime}:=\mathcal{C}_{2}\setminus({\rm col}(P^{2})\cup\{c_{\ell}^{1},c_{\ell}^{2},c_{\ell}^{3},c_{\ell}^{4}:{\ell}\in[i_{0}-1]\})$. Let
$\mathcal{C}_{2}^{1}:=\{j\in\mathcal{C}_{2}^{\prime}:d_{j}^{+}(x_{i_{0}},B)\geq 4(i_{0}-1)+4+|V(P^{2})|\}\ \text{and}$
$\mathcal{C}_{2}^{2}:=\{j\in\mathcal{C}_{2}^{\prime}:d_{j}^{-}(x_{i_{0}},B)\geq 4(i_{0}-1)+4+|V(P^{2})|\}.$
$d_{j}^{+}(x_{i_{0}},A_{j})$ $\geq d_{j}^{+}(x_{i_{0}},A)-|A\setminus A_{1}|-|A_{1}\Delta A_{j}|-|X_{A}\setminus X_{A}^{2}|-|X^{\prime}|$
$\geq\delta^{0}(D_{j})-d_{j}^{+}(x_{i_{0}},B)-|A\setminus A_{1}|-|A_{1}\Delta A_{j}|-|X_{A}\setminus X_{A}^{2}|-|X^{\prime}|$
$\geq\frac{n}{2}-\left(4(i_{0}-1)+4+|V(P^{2})|\right)-2{\epsilon}n-2\delta n-4\delta n-\sqrt{\epsilon}n$
$\geq\left(\frac{1}{2}-3\sqrt{\delta}\right)n.$
Furthermore, $|\mathcal{C}_{2}^{\prime}\setminus\mathcal{C}_{2}^{1}|\geq(1-3\sqrt{\delta})|\mathcal{C}_{2}|$. Notice that $x_{i_{0}}\notin X^{\prime}$. Then, for at least $(1-3\sqrt{\delta})|\mathcal{C}_{2}|$ colors $j\in\mathcal{C}_{2}$, $d_{j}^{-}(x_{i_{0}},B_{j})<\frac{5}{2}\sqrt{\delta}n$, and consequently, $d_{j}^{-}(x_{i_{0}},A_{j})\geq(\frac{1}{2}-3\sqrt{\delta})n$ for the same set of colors. By (C3), we know that $x_{i_{0}}\in A_{j}\cup B_{j}$ for all but at most $6\sqrt{\delta}|\mathcal{C}_{2}|$ colors $j\in\mathcal{C}_{2}$, which leads to $x_{i_{0}}\not\in X$, a contradiction. ∎
By Claim 5.4, for each $x_{i}\in(X\setminus X^{\prime})\cap Y$, there exists a rainbow directed path $Q_{x_{i}}=x_{i}^{1}x_{i}x_{i}^{2}$ with ${\rm col}(Q_{x_{i}})=\{c_{i}^{1},c_{i}^{2}\}\subseteq\mathcal{C}_{2}$ and $x_{i}^{1},x_{i}^{2}\in Z\setminus V_{\rm bad}$. Moreover, each color or endpoint in $Q_{x_{i}}$ can be replaced by an unused color or vertex (if needed). Using Claim 5.3 (i)-(ii), one may connect $P^{2}$ and all rainbow directed paths in $\{Q_{x_{i}}:x_{i}\in X\setminus X^{\prime}\}$ into a single rainbow directed path $P^{3}$ with endpoints $u_{1}\in A\setminus V_{\rm bad}$ and $v_{3}\in B\setminus V_{\rm bad}$, whose length is at most $1027\delta n+2\sqrt{\delta}n+5\sqrt{\epsilon}n+$ 10 .
@@
Choose a maximum rainbow matching in $\{D_{i}^{\pm}[A\setminus V(P^{3}),B\setminus V(P^{3})]:i\in\mathcal{C}_{\rm bad}\setminus{\rm col}(P^{3})\}$, say $\tilde{M}$. If there exists an edge with color $j$ in $\tilde{M}$ such that $D_{j}^{\pm}[A\setminus V(P^{3}\cup\tilde{M}),B\setminus V(P^{3}\cup\tilde{M})]$ contains no $2$-matching, then delete all such edges from $\tilde{M}$ and denote the resulting rainbow directed matching by $M$. Connect $P^{3}$ and all rainbow directed edges in $M$ by Claim 5.3 (i), we obtain a rainbow directed path $P^{4}$ with endpoints $u_{1}\in A\setminus V_{\rm bad}$ and $v_{4}\in B\setminus V_{\rm bad}$, whose length is at most $|E(P^{3})|+16\sqrt{\delta}n\leq 18\sqrt{\delta}n+5\sqrt{\epsilon}n+27\delta n+10$.
@@
By the construction of $P^{4}$, we know that if $w\in V(P^{4})$ is a vertex not in $(V(P^{1})\cap B)\cup V_{\rm bad}\cup X$, then it is possible to avoid $w$ when constructing $P^{4}$; if $i\in{\rm col}(P^{4})$ is a color not in ${\rm col}(P^{1}-A)\cup{\rm col}(M)$, then it is possible to avoid $i$ when constructing $P^{4}$. That is, we can assume that $w\not\in V(P^{4})$ unless $w\in(V(P^{1})\cap B)\cup V_{\rm bad}\cup X^{\prime}$, and $i\not\in{\rm col}(P^{4})$ unless $i\in{\rm col}(P^{1}-A)\cup{\rm col}(M)$.
Furthermore, in the process of choosing rainbow directed paths inside $\mathcal{D}[B]$ (in Step 1), one may use colors from $\mathcal{C}_{1}\cup\mathcal{C}_{\rm bad}^{\prime}$ before other colors and avoid vertices in $X^{\prime}\cup V_{\rm bad}\cup V(M)$. This implies that either $\mathcal{C}_{1}\cup\mathcal{C}_{\rm bad}^{\prime}\subseteq{\rm col}(P^{1})$ or ${\rm col}(P^{1}-A)\subseteq\mathcal{C}_{1}\cup\mathcal{C}_{\rm bad}^{\prime}$. Denote $\tilde{\mathcal{C}}:=(\mathcal{C}_{1}\cup\mathcal{C}_{\rm bad}^{\prime})\setminus{\rm col}(P^{1})$.
Notice that for each color $i\in\mathcal{C}_{2}\setminus{\rm col}(P^{4})$, every vertex $v\in A\setminus V(P^{4})$ has at least $(\frac{1}{2}-3\sqrt{\delta})n-|V(P^{4})|-|B\Delta B_{i}|$ out-neighbors (resp. in-neighbors) in $B\setminus V(P^{4})$ in $D_{i}$. Hence,
@@
Next, we give an application of the transversal blow-up lemma (see [9]) for embedding transversal undirected Hamilton paths inside very dense bipartite graph collections, which can be proved by minor modifications to the proof of [10, Lemma 6.1], and we omit the proof here.
@@
Assume $W,Z\in\{A,B\}$. Let $W^{*}\subseteq W\setminus V_{\rm bad}$ and $Z^{*}\subseteq Z\setminus V_{\rm bad}$, where $|W^{*}|,|Z^{*}|\geq\eta n$, $W^{*}\cap Z^{*}=\emptyset$ and $|W^{*}|-|Z^{*}|=t\in\{0,1\}$. Let $T^{*}=Z^{*}$ if $t=0$ and $T^{*}=W^{*}$ if $t=1$. Let $\mathcal{C^{*}}\subseteq\mathcal{C}$ satisfy $|\mathcal{C^{*}}|=|W^{*}|+|Z^{*}|-1$, where $\mathcal{C^{*}}\subseteq\mathcal{C}_{1}$ if $W=Z$ and $\mathcal{C^{*}}\subseteq\mathcal{C}_{2}$ if $W\neq Z$. Let $W^{-}\subseteq W^{*}$ and $T^{+}\subseteq T^{*}$ with $|W^{-}|,|T^{+}|\geq\frac{\eta n}{8}$. Then there is a transversal undirected Hamilton path in $\{G_{i}[W^{*},Z^{*}]:i\in\mathcal{C}^{*}\}$ starting at $W^{-}$ and ending at $T^{+}$.
Claim 5.6.
Let $Q=x_{1}\ldots x_{s}$ be a rainbow directed path inside $\mathcal{D}$ with $s\leq 19\sqrt{\delta}n$, $V_{\rm bad}\subseteq V(Q)$, $(\mathcal{C}_{1}\cup\mathcal{C}_{\rm bad})\setminus\tilde{\mathcal{C}}\subseteq{\rm col}(Q)$, $x_{1}\in A\setminus V_{\rm bad}$ and $x_{s}\in B\setminus V_{\rm bad}$. If $|\tilde{\mathcal{C}}|-||B\setminus V(Q)|-|A\setminus V(Q)||$ is a nonnegative even integer, then $\mathcal{D}$ has a transversal directed Hamilton cycle.
@@
Using colors in $\tilde{\mathcal{C}}$, one may greedily choose two disjoint rainbow directed paths $P^{5}$ and $P^{6}$ in $\mathcal{D}[A\setminus V(Q)]$ and $\mathcal{D}[B\setminus V(Q)]$ with lengths $\frac{|\tilde{\mathcal{C}}|-(|B\setminus V(Q)|-|A\setminus V(Q)|)}{2}$ and $\frac{|\tilde{\mathcal{C}}|+(|B\setminus V(Q)|-|A\setminus V(Q)|)}{2}$ respectively, whose endpoints are $u_{5},v_{5}\in A\setminus V_{\rm bad}$ and $u_{6},v_{6}\in B\setminus V_{\rm bad}$. (Note that if a rainbow path has length 00, then it contains one vertex.) Applying Claim 5.3 (i)-(ii) to connect $Q,P^{5}$ and $P^{6}$ in turn, we get a rainbow directed path $P$ with endpoints $x_{1}\in A\setminus V_{\rm bad}$ and $v_{6}\in B\setminus V_{\rm bad}$. Clearly, $|V(P)|\leq 19\sqrt{\delta}n+4\sqrt{\delta}n+\eta n+4$ and $|A\setminus V(P)|=|B\setminus V(P)|$.
@@
$|W^{-}|$ $\geq d_{G_{i_{2}}}(v_{6},A)-|V(P)|\geq d_{G_{i_{2}}}(v_{6},A_{i_{2}})-|A_{1}\Delta A_{i_{2}}|-|X_{A}\setminus X_{A}^{2}|-|X^{\prime}|-|V(P)|$
@@
and similarly $|T^{+}|>\left(\frac{1}{2}-2\eta\right)n$. Applying Claim 5.5 yields that $\{G_{i}[W^{*},T^{*}]:i\in\mathcal{C}^{*}\}$ contains a transversal undirected path $P^{\prime}$ starting at $v^{\prime}\in W^{-}$ and ending at $u^{\prime}\in T^{+}$. Thus, $x_{1}Pv_{6}v^{\prime}P^{\prime}u^{\prime}x_{1}$ is a transversal directed Hamilton cycle inside $\mathcal{D}$. ∎
If $n$ and $|\tilde{\mathcal{C}}|$ have the same parity, then in view of Claim 5.6, there exists a transversal directed Hamilton cycle inside $\mathcal{D}$, a contradiction. Hence, it suffices to consider that $n$ and $|\tilde{\mathcal{C}}|$ have the different parity. Therefore, either $|\tilde{\mathcal{C}}|-(|B\setminus V(P^{4})|-|A\setminus V(P^{4})|)$ is a positive odd integer or $|\tilde{\mathcal{C}}|=0$ and $n$ is odd. Based on the definition of $X^{\prime}$, move vertices in $X^{\prime}$ to $Y$ if it is covered by a rainbow directed path from $Z$ to $Y$ (in Step 2).
@@
(i) Suppose that there exists an $i_{0}\in\mathcal{C}\setminus{\rm col}(P^{1}-A)$ such that $D_{i_{0}}[B\setminus(X^{\prime}\cup V(P^{1}))]\neq\emptyset$. Choose ${w_{1}w_{2}}\in E(D_{i_{0}}[B\setminus(X^{\prime}\cup V(P^{1}))])$. Recall that each of $w_{1}$ and $w_{2}$ can be chosen outside $V(P^{4})$ unless it is in $(V(P^{1})\cap B)\cup V_{\rm bad}\cup X^{\prime}$, and $i_{0}$ can be chosen outside ${\rm col}(P^{4})$ unless $i_{0}\in{\rm col}(M)$. If $i_{0}\in{\rm col}(M)$, then let $\tilde{P}^{4}$ be a rainbow directed path obtained by deleting the edge with color $i_{0}$ from $P^{4}$ and connecting its two endpoints by Claim 5.3 (i). It is routine to check that $|B\setminus V(\tilde{P}^{4})|-|A\setminus V(\tilde{P}^{4})|=|B\setminus V({P}^{4})|-|A\setminus V({P}^{4})|$. Let $P^{4}:=\tilde{P}^{4}$ if $i_{0}\in{\rm col}(M)$. Therefore, it suffices to consider $i_{0}\not\in{\rm col}(P^{4})$.
@@
$\bullet$ $w_{1},w_{2}\not\in V_{\rm bad}$. Then $w_{1},w_{2}\not\in V(P^{4})$. Connect $w_{1}$ with the endpoint $v_{4}$ of $P^{4}$ by Claim 5.3 (ii). Denote the resulting rainbow directed path by $Q_{1}$.
@@
$\bullet$ $w_{1}\not\in V_{\rm bad}$ and $w_{2}\in V_{\rm bad}$. Hence $w_{1}\not\in V(P^{4})$ and there exists a rainbow subpath $u_{1}P^{4}w_{2}^{1}w_{2}^{2}w_{2}$ of $P^{4}$ with $w_{2}^{1}\in B\setminus(V_{\rm bad}\cup X^{\prime})$, $w_{2}^{2}\in A\setminus(V_{\rm bad}\cup X^{\prime})$. Replace $w_{2}^{1}w_{2}^{2}w_{2}$ with $w_{1}w_{2}$ in $P^{4}$, and connect $w_{1}$ with the vertex in $N_{P^{4}}^{-}(w_{2}^{1})$ by Claim 5.3 (i). Let $Q_{2}$ be the resulting rainbow directed path. The case for $w_{1}\in V_{\rm bad}$ and $w_{2}\not\in V_{\rm bad}$ is similar.
@@
$\bullet$ $w_{1},w_{2}\in V_{\rm bad}$. By rearranging, one may assume that $w_{1}$ and $w_{2}$ is connected by a rainbow directed path $w_{1}w_{1}^{1}w_{2}^{1}w_{2}^{2}w_{2}$ in $P^{4}$, where $w_{1}^{1},w_{2}^{2}\in A\setminus(V_{\rm bad}\cup X^{\prime})$ and $w_{2}^{1}\in B\setminus(V_{\rm bad}\cup X^{\prime})$. Replace $w_{1}w_{1}^{1}w_{2}^{1}w_{2}^{2}w_{2}$ with $w_{1}w_{2}$ in $P^{4}$. Denote the resulting rainbow directed path by $Q_{3}$.
@@
In each of the above cases, we get a rainbow directed path $Q_{i}$ ($i\in[3]$) such that $|\tilde{\mathcal{C}}|-||B\setminus V(Q_{i})|-|A\setminus V(Q_{i})||$ is a nonnegative even integer. It follows from Claim 5.6 that $\mathcal{D}$ contains a transversal directed Hamilton cycle, a contradiction.
@@
(ii) We will only prove that $D_{i}[A\setminus(V_{\rm bad}\cup X^{\prime}),B]=\emptyset$ for all $i\in\tilde{\mathcal{C}}$, the second statement can be proved by a similar discussion. Suppose that $D_{i_{1}}[A\setminus(V_{\rm bad}\cup X^{\prime}),B]\neq\emptyset$ for some $i_{1}\in\tilde{\mathcal{C}}$. Choose ${z_{1}z_{2}}\in E(D_{i_{1}}[A\setminus(V_{\rm bad}\cup X^{\prime}),B])$. Hence $z_{1}\notin V(P^{4})$.
@@
$\bullet$ $z_{2}\not\in(V(P^{1})\cap B)\cup V_{\rm bad}\cup X^{\prime}$. Then $z_{2}\not\in V(P^{4})$. Connect $z_{2}$ with the endpoint $u_{1}$ of $P^{4}$ by Claim 5.3 (i). Let $Q_{1}$ be the resulting rainbow directed path, and $\tilde{\mathcal{C}}_{1}:=\tilde{\mathcal{C}}\setminus\{i_{1}\}$.
@@
$\bullet$ $z_{2}\in V(P^{1})\cap B$. Then $P^{4}$ can be written as $u_{1}P^{4}z_{2}^{\prime}z_{2}P^{4}v_{4}$, where $z_{2}^{\prime}z_{2}\in E(P^{1})\cap E(D_{i_{2}})$ for some $i_{2}\in\mathcal{C}$. Clearly, $i_{2}\in\mathcal{C}_{2}$ if $z_{2}^{\prime}\in A$, and $i_{2}\in(\mathcal{C}_{1}\cup\mathcal{C}_{\rm bad}^{\prime})\setminus\tilde{\mathcal{C}}$ if $z_{2}^{\prime}\in B$ (since ${\rm col}(P^{1}-A)\subseteq\mathcal{C}_{1}\cup\mathcal{C}_{\rm bad}^{\prime}$). Notice that $V(P^{1})\cap B\cap V_{\rm bad}=\emptyset$. Hence $z_{2}^{\prime}\not\in V_{\rm bad}$. Connect $u_{1}P^{1}z_{2}^{\prime}$ with $z_{1}z_{2}P^{4}v_{4}$ into a single rainbow directed path by Claim 5.3 (i)-(ii). Denote the resulting rainbow directed path by $Q_{2}$. Let $\tilde{\mathcal{C}}_{2}:=\tilde{\mathcal{C}}\setminus\{i_{1}\}$ if $z_{2}^{\prime}\in A$, and $\tilde{\mathcal{C}}_{2}:=(\tilde{\mathcal{C}}\setminus\{i_{1}\})\cup\{i_{2}\}$ if $z_{2}^{\prime}\in B$.
@@
$\bullet$ $z_{2}\in V_{\rm bad}\cup X^{\prime}$. Then $P^{4}$ can be written as $u_{1}P^{4}z_{2}^{1}z_{2}P^{4}v_{4}$, where $z_{2}^{1}\in A\setminus V_{\rm bad}$. Connect $u_{1}P^{4}z_{2}^{1}$ with $z_{1}z_{2}z_{2}^{2}P^{4}v_{4}$ into a single rainbow directed path by Claim 5.3 (ii). Let $Q_{3}$ be the resulting rainbow directed path and $\tilde{\mathcal{C}}_{3}:=\tilde{\mathcal{C}}\setminus\{i_{1}\}$.
@@
In each of the above cases, we get a rainbow directed path $Q_{i}$ ($i\in[3]$) such that $|\tilde{\mathcal{C}}|-||B\setminus V(Q_{i})|-|A\setminus V(Q_{i})||$ is a nonnegative even integer. It follows from Claim 5.6 that $\mathcal{D}$ contains a transversal directed Hamilton cycle, a contradiction. ∎

### v2

@@ 1. $(\dagger)$
For vertices in $X\setminus X^{\prime}$, we consider the following claim. Notice that the position (in $A$ or $B$) of vertices in $X\setminus X^{\prime}$ does not affect our proof, so if $v\in X$ is a vertex moved from $B$ to $A$, then we only need consider $v\in A$. By Lemma 6.1, for each $x_{i}\in(X\setminus X^{\prime})\cap Y$, there exists a rainbow directed path $Q_{x_{i}}=x_{i}^{1}x_{i}x_{i}^{2}$ with ${\rm col}(Q_{x_{i}})=\{c_{i}^{1},c_{i}^{2}\}\subseteq\mathcal{C}_{2}$ and $x_{i}^{1},x_{i}^{2}\in Z\setminus V_{\rm bad}$. Moreover, each color or endpoint in $Q_{x_{i}}$ can be replaced by an unused color or vertex (if needed). Using Lemma 6.3 (i)-(ii), one may connect $P^{2}$ and all rainbow directed paths in $\{Q_{x_{i}}:x_{i}\in X\setminus X^{\prime}\}$ into a single rainbow directed path $P^{3}$ with endpoints $u_{1}\in A\setminus V_{\rm bad}$ and $v_{3}\in B\setminus V_{\rm bad}$, whose length is at most $27\delta n+2\sqrt{\delta}n+5\sqrt{\epsilon}n+10.$
@@
[…]
@@
By the construction of $P^{4}$, we know that if $w\in V(P^{4})$ is a vertex not in $(V(P^{1})\cap B)\cup V_{\rm bad}\cup X$, then it is possible to avoid $w$ when constructing $P^{4}$; if $i\in{\rm col}(P^{4})$ is a color not in ${\rm col}(P^{1}-A)\cup{\rm col}(M)$, then it is possible to avoid $i$ when constructing $P^{4}$. In other words, we may assume that $w\not\in V(P^{4})$ unless $w\in(V(P^{1})\cap B)\cup V_{\rm bad}\cup X^{\prime}$, and $i\not\in{\rm col}(P^{4})$ unless $i\in{\rm col}(P^{1}-A)\cup{\rm col}(M)$.
Furthermore, while selecting rainbow directed paths inside $\mathcal{D}[B]$ (in Step 1), one may give priority to using colors from $\mathcal{C}_{1}\cup\mathcal{C}_{\rm bad}^{\prime}$ and avoid vertices in $X^{\prime}\cup V_{\rm bad}\cup V(M)$. This implies that either $\mathcal{C}_{1}\cup\mathcal{C}_{\rm bad}^{\prime}\subseteq{\rm col}(P^{1})$ or ${\rm col}(P^{1}-A)\subseteq\mathcal{C}_{1}\cup\mathcal{C}_{\rm bad}^{\prime}$. Denote $\tilde{\mathcal{C}}:=(\mathcal{C}_{1}\cup\mathcal{C}_{\rm bad}^{\prime})\setminus{\rm col}(P^{1})$.
Notice that for each color $i\in\mathcal{C}_{2}\setminus{\rm col}(P^{4})$, every vertex $v\in A\setminus V(P^{4})$ has at least $(\frac{1}{2}-3\sqrt{\delta})n-|V(P^{4})|-|B\triangle B_{i}|$ out-neighbors (resp. in-neighbors) in $B\setminus V(P^{4})$ in $D_{i}$. Hence,
@@
The next claim employs Lemma 6.2 to construct a transversal directed Hamilton cycle.
@@
Let $Q=x_{1}\ldots x_{s}$ be a rainbow directed path inside $\mathcal{D}$ with $s\leq 19\sqrt{\delta}n$, $V_{\rm bad}\subseteq V(Q)$, $(\mathcal{C}_{1}\cup\mathcal{C}_{\rm bad})\setminus\tilde{\mathcal{C}}\subseteq{\rm col}(Q)$, $x_{1}\in A\setminus V_{\rm bad}$ and $x_{s}\in B\setminus V_{\rm bad}$. If $|\tilde{\mathcal{C}}|-\big||B\setminus V(Q)|-|A\setminus V(Q)|\big|$ is a nonnegative even integer, then $\mathcal{D}$ has a transversal directed Hamilton cycle.
@@
Using colors in $\tilde{\mathcal{C}}$, one may greedily choose two disjoint rainbow directed paths $P^{5}$ and $P^{6}$ in $\mathcal{D}[A\setminus V(Q)]$ and $\mathcal{D}[B\setminus V(Q)]$ with lengths
$\frac{|\tilde{\mathcal{C}}|-(|B\setminus V(Q)|-|A\setminus V(Q)|)}{2}\ \ \text{and}\ \ \frac{|\tilde{\mathcal{C}}|+(|B\setminus V(Q)|-|A\setminus V(Q)|)}{2}$
respectively, whose endpoints are $u_{5},v_{5}\in A\setminus V_{\rm bad}$ and $u_{6},v_{6}\in B\setminus V_{\rm bad}$. (Note that if a rainbow path has length 0, then it contains one vertex.) Applying Lemma 6.3 (i)-(ii) to connect $Q,P^{5}$ and $P^{6}$ in turn, we get a rainbow directed path $P$ with endpoints $x_{1}\in A\setminus V_{\rm bad}$ and $v_{6}\in B\setminus V_{\rm bad}$. Clearly, $|V(P)|\leq 19\sqrt{\delta}n+4\sqrt{\delta}n+\eta n+4$ and $|A\setminus V(P)|=|B\setminus V(P)|$.
@@
$|W^{-}|$ $\geq d_{G_{i_{2}}}(v_{6},A)-|V(P)|\geq d_{G_{i_{2}}}(v_{6},A_{i_{2}})-|A_{1}\triangle A_{i_{2}}|-|X_{A}\setminus X_{A}^{2}|-|X^{\prime}|-|V(P)|$
@@
and similarly $|T^{+}|>\left(\frac{1}{2}-2\eta\right)n$. Applying Lemma 6.2 yields that $\{G_{i}[W^{*},T^{*}]:i\in\mathcal{C}^{*}\}$ contains a transversal undirected path $P^{\prime}$ starting at $v^{\prime}\in W^{-}$ and ending at $u^{\prime}\in T^{+}$. Thus, $x_{1}Pv_{6}v^{\prime}P^{\prime}u^{\prime}x_{1}$ is a transversal directed Hamilton cycle inside $\mathcal{D}$. ∎
[…]
Based on the definition of $X^{\prime}$, we move vertices in $X^{\prime}$ to $Y$ if it is covered by a rainbow directed path from $Z$ to $Y$ (in Step 2). The following claim considers the local structure of $D_{i}$.
@@
(i) Suppose that there exists an $i_{0}\in\mathcal{C}\setminus{\rm col}(P^{1}-A)$ such that $D_{i_{0}}[B\setminus(X^{\prime}\cup V(P^{1}))]\neq\emptyset$. Choose ${w_{1}w_{2}}\in E(D_{i_{0}}[B\setminus(X^{\prime}\cup V(P^{1}))])$. Recall that each of $w_{1}$ and $w_{2}$ can be chosen outside $V(P^{4})$ unless it is in $(V(P^{1})\cap B)\cup V_{\rm bad}\cup X^{\prime}$, and $i_{0}$ can be chosen outside ${\rm col}(P^{4})$ unless $i_{0}\in{\rm col}(M)$. If $i_{0}\in{\rm col}(M)$, then let $\tilde{P}^{4}$ be a rainbow directed path obtained by deleting the edge with color $i_{0}$ from $P^{4}$ and connecting its two endpoints by Lemma 6.3 (i). It is routine to check that $|B\setminus V(\tilde{P}^{4})|-|A\setminus V(\tilde{P}^{4})|=|B\setminus V({P}^{4})|-|A\setminus V({P}^{4})|$. Let $P^{4}:=\tilde{P}^{4}$ if $i_{0}\in{\rm col}(M)$, which reduces our proof to the case where $i_{0}\not\in{\rm col}(P^{4})$. We consider the following several cases.
@@
$\bullet$ Suppose $w_{1},w_{2}\not\in V_{\rm bad}$. Then $w_{1},w_{2}\not\in V(P^{4})$. Connect $w_{1}$ with the endpoint $v_{4}$ of $P^{4}$ by Lemma 6.3 (ii). Denote the resulting rainbow directed path by $Q_{1}$.
@@
$\bullet$ Suppose $w_{1}\not\in V_{\rm bad}$ and $w_{2}\in V_{\rm bad}$. Hence $w_{1}\not\in V(P^{4})$ and there exists a rainbow subpath $u_{1}P^{4}w_{2}^{1}w_{2}^{2}w_{2}$ of $P^{4}$ with $w_{2}^{1}\in B\setminus(V_{\rm bad}\cup X^{\prime})$, $w_{2}^{2}\in A\setminus(V_{\rm bad}\cup X^{\prime})$. Replace $w_{2}^{1}w_{2}^{2}w_{2}$ with $w_{1}w_{2}$ in $P^{4}$, and connect $w_{1}$ with the vertex in $N_{P^{4}}^{-}(w_{2}^{1})$ by Lemma 6.3 (i). Let $Q_{2}$ be the resulting rainbow directed path. The case for $w_{1}\in V_{\rm bad}$ and $w_{2}\not\in V_{\rm bad}$ is similar.
@@
$\bullet$ Suppose $w_{1},w_{2}\in V_{\rm bad}$. By rearranging, one may assume that $w_{1}$ and $w_{2}$ are connected by a rainbow directed path $w_{1}w_{1}^{1}w_{2}^{1}w_{2}^{2}w_{2}$ in $P^{4}$, where $w_{1}^{1},w_{2}^{2}\in A\setminus(V_{\rm bad}\cup X^{\prime})$ and $w_{2}^{1}\in B\setminus(V_{\rm bad}\cup X^{\prime})$. Replace $w_{1}w_{1}^{1}w_{2}^{1}w_{2}^{2}w_{2}$ with $w_{1}w_{2}$ in $P^{4}$; and denote the resulting rainbow directed path by $Q_{3}$.
@@
In each of the above cases, we get a rainbow directed path $Q_{i}$ ($i\in[3]$) such that $|\tilde{\mathcal{C}}|-\big||B\setminus V(Q_{i})|-|A\setminus V(Q_{i})|\big|$ is a nonnegative even integer. It follows from Claim 6.6 that $\mathcal{D}$ contains a transversal directed Hamilton cycle, a contradiction.
@@
(ii) We will only prove that $D_{i}[A\setminus(V_{\rm bad}\cup X^{\prime}),B]=\emptyset$ for all $i\in\tilde{\mathcal{C}}$, the second statement can be proved by a similar discussion. Suppose that $D_{i_{1}}[A\setminus(V_{\rm bad}\cup X^{\prime}),B]\neq\emptyset$ for some $i_{1}\in\tilde{\mathcal{C}}$. Choose ${z_{1}z_{2}}\in E(D_{i_{1}}[A\setminus(V_{\rm bad}\cup X^{\prime}),B])$. Hence $z_{1}\notin V(P^{4})$. Next, we consider several cases.
@@
$\bullet$ Suppose $z_{2}\not\in(V(P^{1})\cap B)\cup V_{\rm bad}\cup X^{\prime}$. Then $z_{2}\not\in V(P^{4})$. Using Lemma 6.3 (i), connect $z_{2}$ to the endpoint $u_{1}$ of $P^{4}$. Denote the resulting rainbow directed path by $Q_{1}$, and set $\tilde{\mathcal{C}}_{1}:=\tilde{\mathcal{C}}\setminus\{i_{1}\}$.
@@
$\bullet$ Suppose $z_{2}\in V(P^{1})\cap B$. Then $P^{4}$ can be written as $u_{1}P^{4}z_{2}^{\prime}z_{2}P^{4}v_{4}$, where $z_{2}^{\prime}z_{2}\in E(P^{1})\cap E(D_{i_{2}})$ for some $i_{2}\in\mathcal{C}$. Clearly, $i_{2}\in\mathcal{C}_{2}$ if $z_{2}^{\prime}\in A$, and $i_{2}\in(\mathcal{C}_{1}\cup\mathcal{C}_{\rm bad}^{\prime})\setminus\tilde{\mathcal{C}}$ if $z_{2}^{\prime}\in B$ (since ${\rm col}(P^{1}-A)\subseteq\mathcal{C}_{1}\cup\mathcal{C}_{\rm bad}^{\prime}$). Notice that $V(P^{1})\cap B\cap V_{\rm bad}=\emptyset$. Hence $z_{2}^{\prime}\not\in V_{\rm bad}$. Apply Lemma 6.3 (i)-(ii) to connect $u_{1}P^{1}z_{2}^{\prime}$ and $z_{1}z_{2}P^{4}v_{4}$ into a single rainbow directed path, which we denote by $Q_{2}$. Let $\tilde{\mathcal{C}}_{2}:=\tilde{\mathcal{C}}\setminus\{i_{1}\}$ if $z_{2}^{\prime}\in A$, and $\tilde{\mathcal{C}}_{2}:=(\tilde{\mathcal{C}}\setminus\{i_{1}\})\cup\{i_{2}\}$ if $z_{2}^{\prime}\in B$.
@@
$\bullet$ Suppose $z_{2}\in V_{\rm bad}\cup X^{\prime}$. Then $P^{4}$ can be written as $u_{1}P^{4}z_{2}^{1}z_{2}P^{4}v_{4}$, where $z_{2}^{1}\in A\setminus V_{\rm bad}$. By Lemma 6.3 (ii), we connect $u_{1}P^{4}z_{2}^{1}$ with $z_{1}z_{2}z_{2}^{2}P^{4}v_{4}$ into a single rainbow directed path, say $Q_{3}$. Set $\tilde{\mathcal{C}}_{3}:=\tilde{\mathcal{C}}\setminus\{i_{1}\}$.
@@
In each of the above cases, we get a rainbow directed path $Q_{i}$ ($i\in[3]$) such that $|\tilde{\mathcal{C}}|-\big||B\setminus V(Q_{i})|-|A\setminus V(Q_{i})|\big|$ is a nonnegative even integer. It follows from Claim 6.6 that $\mathcal{D}$ contains a transversal directed Hamilton cycle, a contradiction. ∎